rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#14655 — bhs-3b-a9

Attached to Project— Network
Incident
Beauharnois
CLOSED
100%
We loss access to bhs-3b-a9. We investiguate.
Date:  Thursday, 01 June 2017, 14:45PM
Reason for closing:  Done
Comment by OVH - Thursday, 01 June 2017, 07:44AM

The equipement come back and then down again.
We contact Cisco in emergency.

For now, service is stable.

We investiguate with cisco.


Comment by OVH - Thursday, 01 June 2017, 07:52AM

Cisco investiguate about this log:

2017-06-01T06:29:10+02:00 198.27.73.96 bhs-3b-a9 3307013 ba [local7.crit] === 3307013: LC/0/1/CPU0:Jun 1 06:29:10 CEST: fib_mgr[183]: %ROUTING-FIB-2-OOR : FIB has run out of DATA_TYPE_TABLE_SET resource memory. No more route updates will be handled by FIB. Please delete routes, and then clear FIB on this node to resume normal operation.
2017-06-01T06:31:17+02:00 198.27.73.96 bhs-3b-a9 3307026 ba [local7.crit] === 3307026: LC/0/0/CPU0:Jun 1 06:31:17 CEST: fib_mgr[183]: %ROUTING-FIB-2-OOR : FIB has run out of DATA_TYPE_TABLE_SET resource memory. No more route updates will be handled by FIB. Please delete routes, and then clear FIB on this node to resume normal operation.
2017-06-01T06:41:40+02:00 198.27.73.96 bhs-3b-a9 3307127 be [local7.info] === 3307127: LC/0/1/CPU0:Jun 1 06:41:40 CEST: fib_mgr[183]: %ROUTING-FIB-6-RSRC_OK : FIB resource state has returned to normal. FIB has exited resource constrained operation and normal forwarding has been restored
2017-06-01T06:41:47+02:00 198.27.73.96 bhs-3b-a9 3307129 be [local7.info] === 3307129: LC/0/0/CPU0:Jun 1 06:41:47 CEST: fib_mgr[183]: %ROUTING-FIB-6-RSRC_OK : FIB resource state has returned to normal. FIB has exited resource constrained operation and normal forwarding has been restored


Comment by OVH - Thursday, 01 June 2017, 10:30AM

Cisco talks about a possible known bug.
We will continue the investigation with Cisco.
All IP returned at 07:02 GMT+2 (servers IP) and 07:24 (IPFO)

We will upgrade the couple as soon as possible.


Comment by OVH - Thursday, 01 June 2017, 12:40PM

During the incident of this morning, nous moved some vlans by shutting them on B side to restaure the service quickly. All the traffic for these vlans passes through A side.
Here are the vlans:
290, 292, 294 to 303, 305 to 310


Comment by OVH - Thursday, 01 June 2017, 13:19PM

We see again some loss.
We investiguate.


Comment by OVH - Thursday, 01 June 2017, 14:44PM

Traffic is now stable since 13:41 GMT+2